# The Nuclear Star Cluster of M 74: a fossil record of the very early stages of a star-forming galaxy

### v1

@@ 3.1 The nuclear star cluster (NSC) of M 74
The NSC in M 74 was first identified with HST by Ganda2009 and later morphologically analyzed by Georgiev2014 and Georgiev2016. A detailed photometric study by Hoyer2023b, combining JWST and HST imaging, confirmed these results and revealed additional structural complexity. They derived a NSC effective radius of $R_{\rm eff}{\sim}5$pc, an ellipticity ($\epsilon$) of 0.05, a Sérsic index between 2 and 3, using ultraviolet, optical and near-infrared band filters. JWST Mid-Infrared Instrument (MIRI) data gave a larger size of $R_{\rm eff}{\sim}12$ pc, higher ellipticity ($\epsilon{\sim}0.4$), and a Sérsic profile close to exponential ($n{\sim}1.5$). This suggests multiple components or wavelength-dependent substructures. Hoyer2023b discussed different explanations for the mid-infrared emission, including the contribution from a weak AGN, dust released by asymptotic-giant-branch (AGB) stars, an infalling star cluster, or a circumnuclear disk.
Spectral Energy Distribution (SED) fitting across ten bands from the ultraviolet to the near infrared yielded an old age of $8\pm 3$ Gyr and a subsolar metallicity $Z=0.012\pm 0.006$. This suggested that the NSC was not involved in relatively recent star formation (within few Gyr). The inclusion of MIRI data degraded the fit. The PHANGS-JWST observations showed that the NSC resides in the central cavity devoid of gas and dust, contrasting with the dense and complex web of filaments of the galaxy (see Sect. 3). The old ages of the NSC suggest that the cavity has persisted for several Gyr at least. It may have formed just by gas depletion combined with some mechanism preventing gas from inflowing to the center, or by the evacuation of the previously present gas. Hoyer2023b obtained a NSC stellar mass of ${\sim}1.26\times 10^{7}$ M. While this is a relatively massive NSC compared to the average (e.g., Fahrion2022b), it is less massive than NSCs in galaxies of similar masses as M 74 (Hoyer2023b). This is consistent with limited subsequent in-situ growth due to the gas-free environment.
Archival Chandra data reveal an X-ray source associated with the NSC (Lehmer2024), whose X-ray spectrum is consistent with one or more low-mass X-ray binaries (LMXBs) in which a low-mass star accretes on a stellar-mass black hole or a neutron star. The presence of LMXBs is a signature of old stellar populations and they are usually found in elliptical galaxies and globular clusters. While an AGN origin cannot be ruled out, it is disfavored given the association of the X-ray source with the NSC.
@@ 4.1 Spectrophotometric decomposition of the NSC: C2D
As mentioned in Sect. 1, it is important to disentangle the NSC light from the other galaxy components, overlapped in the central region along the LOS, to analyze the properties of the NSC more accurately. For this purpose, we used C2D, a multicomponent spectrophotometric decomposition code (MendezAbreu2019). C2D creates one quasi-monochromatic image for each spectral pixel (i.e. per wavelength), and then performs a multicomponent photometric decomposition of those images. The photometric decomposition of the galaxy surface brightness distribution in each image is done in two dimensions with the code GAlaxy Surface Photometry 2 Dimensional Decomposition (GASP2D, MendezAbreu2008). GASP2D adopts a Levenberg-Marquardt algorithm to estimate the structural parameters that fit the two-dimensional surface brightness of the galaxy. In the first version, it was optimized to fit a bulge and a disk (MendezAbreu2008), while it was later adapted to fit also bars (MendezAbreu2014). In this work, we adapted GASP2D to fit an unresolved NSC, in addition to a bulge and a disk (Sect. 4.1.2).
@@ 4.1.1 Preparation of the data cube of the central region of the galaxy
Since this work focuses on the central region of the galaxy, we first cropped the original PHANGS-MUSE mosaic to the central arcmin, and all the analysis in this work was conducted on this cropped data cube. This ensured a successful NSC decomposition, avoiding the need to handle the full mosaic, which is computationally expensive. This also avoided offsets in the stellar population parameters between different MUSE pointings, which were detected in PHANGS-MUSE large mosaics due to biases in the sky subtraction (Emsellem2022). An image obtained from the cropped data cube of the central arcmin, using the full MUSE wavelength range, is shown in Fig. 1. We also created a mask to exclude bright foreground stars from the analysis described below (see, e.g., Fig. 2 and 3).
For the full analysis, we used the wavelength range between 4800 and 5500 $\mathrm{\SIUnitSymbolAngstrom}$. Within the MUSE spectral coverage, this is the wavelength range containing most of the spectral features sensitive to the stellar population parameters of interest (e.g., Pinna2019b; Pinna2019a; Sattler2023; Sattler2025), while not being as affected as the redder range by sky residuals and imperfections in the modeling of molecular bands of the stellar population models. Using this wavelength range, we applied a Voronoi binning (Cappellari2003), using the VORBIN Python package. This method spatially bins the data in an adaptive way, preserving the maximum possible spatial resolution, to reach a target signal-to-noise ratio (S/N) per bin. We initially used a target S/N =25=25 to avoid large uncertainties in the calculation of the kinematic parameters (see below), but at the same time preserving a spatial resolution very close to the one of the unbinned cube, especially in the very central region where our interest resides. Preserving the highest possible spatial resolution is important for the spectro-photometric decomposition, and binning more would have smoothed the spatial information that C2D needs to fit each of the quasi-monochromatic images. We ensured that all included spaxels before binning had a S/N>3>3.
@@
C2D performs the spectro-photometric decomposition in each individual quasi-monochromatic image, which is created with the flux at the same wavelength position for all spaxels. Thus, it is important not to mix in the same image information from different restframe positions in the absorption or emission spectral features. In the observed spectra, spectral features are shifted due to Doppler effects, in an amount that is different according to the gas and stellar kinematics. Therefore, before applying C2D, we need all spaxels to have the same velocity ($V=0$ km s-1 at restframe) and the same line broadening (due to the velocity dispersion). For that, we corrected the data cube from the stellar kinematics of each spaxel. We calculated the LOS velocity and velocity dispersion in each spaxel of our cropped and Voronoi binned data cube, as explained in Sect. 4.2. Accounting for the kinematics in each spaxel, we shifted the entire data cube to the rest frame and we convolved it to the maximum measured velocity dispersion in the analyzed data cube (${\sim}155$ km s-1). Since M 74 is a star-forming galaxy, spaxels in the disk region have emission lines, often strong, in their spectra. We tested C2D on the kinematic corrected data cube both with and without previously subtracting emission lines. The emission cleaning for these tests was conducted using pPXF (Cappellari2004, see also Sect 4.2). We finally decided not to feed C2D with the emission-subtracted data cube, but rather use the cube including the emission. This choice avoided introducing any additional bias when subtracting emission lines from spectra, and gave better C2D fits. Since emission characterizes specific regions, such as the disk-dominated region, and not others like the very central, NSC-dominated region, preserving these differences helped C2D to perform a better multicomponent fit. MendezAbreu2019 showed the capability of C2D to process emission lines properly.
@@ 4.1.2 Application of C2D to M 74
For this work, we adapted C2D and GASP2D to include the following three components: a NSC, a bulge and a disk. This choice was, on the one hand, motivated by the fact that this galaxy was fitted with a bulge and a disk by Salo2015 as part of the S4G sample (Sheth2010). On the other hand, we are mostly interested in the NSC, which is not spatially resolved in MUSE data as its size is smaller than the PSF size (Sect. 3.1). Its light is spread by the PSF in a larger region than its size, and we follow other studies modeling the NSC as a PSF to disentangle its spectra from the galaxy spectra (e.g., Johnston2020; Fahrion2021).
@@
For the bulge and the disk we used respectively a Sérsic and a broken-exponential two-dimensional profiles (see MendezAbreu2017 for the specific implementation). We extensively tested different combinations of free and fixed parameters of these profiles, resulting in a better decomposition with a certain number of them fixed, as described below. Furthermore, the impact of small errors in the bulge and disk parameters on the NSC spectra, which is the goal of the decomposition, is negligible. The effective intensity, the effective radius and the Sérsic parameter of the bulge, as well as the central intensity of the disk, were kept free and fitted by the code, while the ellipticity and the position angle of both the bulge and the disk were fixed to the values from the S4G decomposition by Salo2015, who fitted the disk with a single exponential. We checked that the effective radius and the Sérsic parameter of the bulge resulting from the fit were very similar to the values from Salo2015. We did not expect to find the same values since they used a different wavelength range compared to ours. Before choosing the disk profile as a broken exponential, we performed a preliminary morphological analysis, fitting ellipses to the isophotes, in an image obtained from the entire PHANGS-MUSE mosaic, integrating all wavelengths in the analyzed spectral range. A one-dimensional radial profile, obtained by averaging azimuthally, follows a broken exponential, with the outer-profile scale length matching the disk scale length from S4G (Salo2015, using a single exponential). This is in agreement with the fact that the S4G team performed the morphological fit on a much larger spatial scale. We measured that the inner exponential has a scale length of 17.4 arcsec (${\sim}827$ pc) while the break radius is located at 30 arcsec (${\sim}1.4$ kpc) from the center. While our cropped data cube is dominated by the inner exponential, we fixed both these parameters in the C2D decomposition.
@@
We fitted the NSC with a Moffat function, as this best characterizes the PSF (Sect. 2). We adopted a FWHM of 0.6 arcsec (three MUSE spaxels, Sect. 2), while we tested that the results were not affected by variations of $\pm 0.2$ arcsec (one spaxel). To test this, we repeated the full decomposition and spectral-fitting process with a FWHM of 0.4 and 0.8 arcsec, finding that the differences in the NSC stellar populations were negligible. The FWHM and the power index (2.8, see Sect. 2) of the Moffat function were fixed in the fitting process, while the PSF (NSC) central intensity was kept free and fitted. With our new setup, GASP2D fitted each quasi-monochromatic image with a combination of a bulge, a disk and a NSC. This resulted in three data cubes, one for each component. Images resulting from integrating all wavelengths for each data cube are shown in Fig. 2. The light fractions of each component over the total are shown in two dimensions in Fig 3. This figure shows that the NSC dominates the light (with a light fraction above 60%) in the central spaxels within one PSF FWHM, while 100% of the light was assigned to the disk from a certain radius (${\sim}7$ arcsec) outwards.
@@
Since we are interested in the stellar-population properties of the NSC and its comparison with the surrounding galaxy, we will use for the rest of the analysis a data cube for the NSC and another for the rest of the galaxy. The latter was obtained by combining the bulge and the disk data cubes. We preferred this approach instead of considering separately the bulge and the disk, to minimize the impact of a potential suboptimal decomposition of these two morphological structures. We extracted integrated spectra in an aperture of radius $1\sigma$ of the PSF FWHM (${\sim}0.26$ arcsec $\approx 12$ pc) from these two data cubes and, for comparison, from the original kinematic-corrected data cube before being processed by C2D (Sect. 4.1.1). Taking into account the central region of the PSF ensures that small errors in the PSF estimate and variations of the PSF size with wavelength have no impact on the extracted properties of the NSC (see also previous paragraph). These spectra are shown and compared to each other, for $1\sigma$ PSF, in Fig. 4.
Figure 4: Spectra obtained by integrating, within an aperture of $1\sigma$ of the PSF, the prepared (original) data cube before the spectrophotometric decomposition (blue solid), the NSC data cube (red dashed), and the host-galaxy data cube (green dashed).
@@ 4.2 Spectral fitting
We performed full spectral fitting using the latest Python implementation of the penalized pixel-fitting (pPXF) method (Cappellari2004; Cappellari2017; Cappellari2023). This models an observed spectrum as a linear combination of simple stellar population (SSP) templates, convolved with a line-of-sight velocity distribution (LOSVD) described by a Gauss–Hermite series. The resulting (light- or mass-fraction) weights, assigned to the SSP models to obtain the best-fit spectrum, are used for the stellar-population analysis, while the LOSVD is used to extract the kinematic parameters. Before applying the spectro-photometric decomposition with C2D, we corrected the original data cube from stellar-velocity shift and broadening (Sect. 4.1.1). Stellar velocities and velocity dispersions were derived by fitting the original central data cube, Voronoi binned to a target S/N=25=25, with pPXF using an additive polynomial of degree 8.
@@
We used in this work the semi-empirical sMILES SSP models (Knowles2023), the most up-to-date version of the MILES SSP models (Vazdekis2010; Vazdekis2015). sMILES models are based on BaSTI isochrones (Pietrinferni2004; Pietrinferni2006) and offer five different values of the [Mg/Fe] abundances, covering the range between -0.2 and 0.6 dex (in steps of 0.2 dex). Throughout this paper, we will use [Mg/Fe] as a proxy for the enhancement in $\alpha$ elements. sMILES models cover ages between 0.03 and 14 Gyr, with 53 different values, and ten values of total metallicities [M/H] between -1.79 and 0.26 dex. This gives a total of 2650 SSP models. Since each of them corresponds to one combination of age, [M/H] and [Mg/Fe], coupling these models with pPXF allows us to extract the distribution of these three stellar-population parameters from each spectrum (given by the distribution of weights given to the different SSPs). These models cover a wavelength range between 3540 and 7409.6 $\mathrm{\SIUnitSymbolAngstrom}$, at a spectral resolution of 2.51 $\mathrm{\SIUnitSymbolAngstrom}$ at FWHM. We chose the set of SSPs that adopted a bimodal initial mass function with a logarithmic slope of 1.3 for the segment of massive stars (Vazdekis2003). These models were normalized to correspond to one solar mass each, so that mass-weighted results are obtained from pPXF if they are not renormalized. They need to be renormalized to their own median to obtain light-weighted results. We used in this work both normalizations, to obtain light-weighted stellar population maps and average results, and mass-weighted results to allow us to reconstruct SFHs in terms of mass fractions corresponding to different stellar populations.
@@
Our general approach for the stellar-population analysis was to average results from Monte Carlo simulations of unregularized pPXF fits, following Cappellari2023. We used this approach to fit the three spectra extracted from the PSF aperture (Sect. 4.1.2), i.e. the NSC, the host-galaxy, and for comparison the spectrum from the original integrated data cube. We performed 100 Monte Carlo realizations for each spectrum. For each realization, we perturbed the spectrum with random, wavelength-dependent noise drawn from a Gaussian distribution. The noise level was taken from the residuals at the specific wavelength, from an initial (unregularized) pPXF fit. From the distribution of SSP weights for the 100 realizations, we estimated the mean ages, [M/H] and [Mg/Fe], and their statistical uncertainties (as standard deviations), for each of the three original spectra. This allowed us not to use any regularization on the SSP weights during the pPXF fits.
@@
The resulting host-galaxy data cube from the decomposition was rebinned to a higher S/N of 100 before performing the spectral fitting for the stellar population analysis, ensuring robust results and lower uncertainties in the outer fainter region. This still left us with 6766 Voronoi bins, which made the Monte Carlo approach time-consuming and computationally expensive for the entire data cube. Thus, we decided to map the stellar-population parameters of the host galaxy using pPXF regularized fits with a unit regularization parameter. While one single unregularized fit gives discrete and noisy solutions, some level of regularization ensures smoother and more realistic stellar-population solutions (e.g., Cappellari2017). However, regularization above a certain level can lead to biases in the results and dilute features in the SFHs. Thus, we used the approach presented in Pinna2019a to choose the regularization parameter (described as follows, see also Pinna2019b; Martig2021; Sattler2023; Sattler2025). After rescaling the pPXF noise parameter as the one necessary to obtain a reduced $\chi^{2}=1$ with no regularization, we choose a regularization well below a certain maximum calibration value. We defined the calibration regularization as the minimum value increasing $\chi^{2}$, from the case with no regularization, by $\Delta\chi^{2}\approx\sqrt{2N_{\text{goodpix}}}$, where $N_{\text{goodpix}}$ is the number of fitted (spectral) pixels (McDermid2015; Cappellari2017; Boecker2020). For comparison, we also mapped the stellar populations from the original data cube before the C2D decomposition, using this same approach (S/N=100=100, multiplicative polynomial of degree 8 and unit regularization parameter).

### v2

@@ 3.1 The nuclear star cluster (NSC) of M 74
M 74’s NSC was first identified with HST by Ganda et al. (2009) and later morphologically analyzed by Georgiev and Böker (2014) and Georgiev et al. (2016). A detailed study by Hoyer et al. (2023b), combining JWST and HST imaging, confirmed these results and revealed additional structural complexity. They derived a NSC effective radius of $R_{\rm eff}{\sim}5$pc, an ellipticity ($\epsilon$) of 0.05, a Sérsic index between 2 and 3, using ultraviolet, optical and near-infrared band filters. JWST Mid-Infrared Instrument (MIRI) data gave a larger size of $R_{\rm eff}{\sim}12$ pc, higher ellipticity ($\epsilon{\sim}0.4$), and a Sérsic profile close to exponential ($n{\sim}1.5$). This suggests multiple components or wavelength-dependent substructures. Hoyer et al. (2023b) discussed different explanations for the mid-infrared emission, including the contribution from a weak AGN, dust released by asymptotic-giant-branch (AGB) stars, an infalling star cluster, or a circumnuclear disk.
Spectral Energy Distribution (SED) fitting across ten bands from the ultraviolet to the near infrared yielded an old age of $8\pm 3$ Gyr and a subsolar metallicity $Z=0.012\pm 0.006$. This suggested that the NSC was not involved in relatively recent star formation (within few Gyr). The inclusion of MIRI data degraded the fit. The PHANGS-JWST observations showed that the NSC resides in the central cavity devoid of gas and dust, contrasting with the dense and complex web of filaments of the galaxy (see Sect. 3). The old ages of the NSC suggest that the cavity has persisted for several Gyr at least. It may have formed just by gas depletion combined with some mechanism preventing gas from inflowing to the center, or by the evacuation of the previously present gas. Hoyer et al. (2023b) obtained a NSC stellar mass of ${\sim}1.26\times 10^{7}$ M. While this is a relatively massive NSC compared to the average (e.g., Fahrion et al. 2022a), it is less massive than NSCs in galaxies of similar masses as M 74 (Hoyer et al. 2023b). This is consistent with limited subsequent in-situ growth due to the gas-free environment.
Archival Chandra data reveal an X-ray source associated with the NSC (Lehmer et al. 2024), whose X-ray spectrum is consistent with one or more low-mass X-ray binaries (LMXBs) in which a low-mass star accretes on a stellar-mass black hole or a neutron star. The presence of LMXBs is a signature of old stellar populations and they are usually found in elliptical galaxies and globular clusters. While an AGN origin cannot be ruled out, it is disfavored given the association of the X-ray source with the NSC.
@@ 4 Methods
In this paper, we analyze the stellar populations of the NSC in M 74 in the context of the host galaxy. We first disentangled the NSC light from its host using a spectro-photometric decomposition method, yielding separate data cubes for each component (Sect. 4.1). For the stellar population analysis, we used two overall approaches. One consisted of fitting spectra integrated over a PSF aperture (Sect. 4.1.2), separately for the NSC and the host galaxy (one each, Fig. 4). Secondly, we fitted the spectra of each spaxel in host-galaxy data cube, allowing the spatially resolved analysis of its stellar populations. All the details about the spectral fitting process are given in Sect. 4.2.
@@ 4.1 Spectrophotometric decomposition of the NSC: C2D
As mentioned in Sect. 1, it is important to disentangle the NSC light from the other galaxy components, overlapped in the central region along the LOS, to analyze the properties of the NSC more accurately. For this purpose, we used C2D, a multicomponent spectrophotometric decomposition code (Méndez-Abreu et al. 2019). C2D creates one quasi-monochromatic image for each spectral pixel (i.e. per wavelength), and then performs a multicomponent photometric decomposition of those images. The photometric decomposition of the galaxy surface brightness distribution in each image is done in two dimensions with the code GAlaxy Surface Photometry 2 Dimensional Decomposition (GASP2D, Méndez-Abreu et al. 2008). GASP2D adopts a Levenberg-Marquardt algorithm to estimate the structural parameters that fit the two-dimensional surface brightness of the galaxy. In the first version, it was optimized to fit a bulge and a disk (Méndez-Abreu et al. 2008), while it was later adapted to fit also bars (Méndez-Abreu et al. 2014). In this work, we adapted GASP2D to fit an unresolved NSC, in addition to a bulge and a disk (Sect. 4.1.2).
@@ 4.1.1 Preparation of the data cube of the central region of the galaxy
In this section, we describe all the steps needed to prepare the data for the C2D spectrophotometric decomposition. We used, for the full analysis presented in this paper, the kinematic-corrected data cube resulting from all the steps described here.
  1. 1.
@@
    Cropping the mosaic. Since this work focuses on the central region of the galaxy, we first cropped the original PHANGS-MUSE mosaic to the central arcmin. This ensured a successful NSC decomposition, avoiding the need to handle the full mosaic, which is computationally expensive. This also avoided offsets in the stellar population parameters between different MUSE pointings, which were detected in PHANGS-MUSE large mosaics due to biases in the sky subtraction (Emsellem et al. 2022). An image obtained from the cropped data cube of the central arcmin, using the full MUSE wavelength range, is shown in Fig. 1. We also created a mask to exclude bright foreground stars from the analysis described below (see, e.g., Fig. 2 and 3).
  2. 2.
[…]
  3. 3.
@@
    Correction from stellar kinematics. C2D performs the spectro-photometric decomposition in each individual quasi-monochromatic image, which is created with the flux in the same spectral pixel, i.e. at the same wavelength position, for all spaxels. When extracting the images, the spectral features in the data cube (e.g. absorption or emission lines) must be located in the same spectral pixels (same wavelength) for all spaxels. Otherwise, one image may mix information from different rest-frame wavelengths. In the observed spectra, spectral features are shifted not only an amount corresponding to the redshift of the galaxy, but also due to (gas- and stellar-) kinematic Doppler effects. Moreover, the observed width of these spectral features is affected by the stellar and gas velocity dispersion. Therefore, before applying C2D, we need all spaxels to have the same velocity ($V=0$ km s-1 at restframe) and the same line broadening (velocity dispersion). For that, we corrected the data cube from the stellar kinematics of each spaxel. We calculated the LOS velocity and velocity dispersion in each spaxel of our cropped and Voronoi binned data cube, as explained in Sect. 4.2. Accounting for the kinematics in each spaxel, we shifted the entire data cube to the rest frame and we convolved it to the maximum measured velocity dispersion in the analyzed data cube (${\sim}155$ km s-1).
  4. 4.
    Treatment of the emission lines. Since M 74 is a star-forming galaxy, spaxels in the disk region have emission lines, often strong, in their spectra. We tested C2D on the kinematic corrected data cube both with and without previously subtracting emission lines. The emission cleaning for these tests was conducted using pPXF (Cappellari and Emsellem 2004, see also Sect 4.2). We finally decided not to feed C2D with the emission-subtracted data cube, but rather use the cube including the emission. This choice avoided introducing any additional bias when subtracting emission lines from spectra, and gave better C2D fits. Since emission characterizes specific regions, such as the disk-dominated region, and not others like the very central, NSC-dominated region, preserving these differences helped C2D to perform a better multicomponent fit. Méndez-Abreu et al. (2019) showed the capability of C2D to process emission lines properly.
@@ 4.1.2 Application of C2D to M 74
This work analyzes the stellar populations of M 74’s NSC and compares them with those of the host galaxy. Because the NSC is unresolved, its light was measured within a PSF aperture, while the surrounding galaxy was spatially resolved using the full capabilities of the IFS data. This enables a direct comparison between the NSC and its immediate environment, placing its formation in the context of the host galaxy. We adopt C2D over other spectrophotometric decomposition codes (e.g., BUDDI, Johnston et al. 2017, see also Sect. 1), because it uniquely preserves both spectral and spatial information for each component. This allows us not only to subtract the PSF-weighted NSC light but also map the host-galaxy properties, fully exploiting the spatial information from IFS.
@@
For this work, we adapted C2D and GASP2D to include the following three components: a NSC, a bulge and a disk. This choice was, on the one hand, motivated by the fact that this galaxy was fitted with a bulge and a disk by Salo et al. (2015) as part of the S4G sample (Sheth et al. 2010). On the other hand, we are mostly interested in the NSC, which is not spatially resolved in MUSE data as its size is smaller than the PSF size (Sect. 3.1). Its light is spread by the PSF in a larger region than its size, and we follow other studies modeling the NSC as a PSF to disentangle its spectra from the galaxy spectra (e.g., Johnston et al. 2020; Fahrion et al. 2021). We fitted the NSC using the Moffat function from the PSF modeling in Emsellem et al. (2022, see Sect. 2 for more details). While we adopted a FWHM of 0.6 arcsec (three MUSE spaxels), we tested that the results were not affected by variations of $\pm 0.2$ arcsec (one spaxel). To test this, we repeated the full decomposition and spectral-fitting process with a FWHM of 0.4 and 0.8 arcsec, finding that the differences in the NSC stellar populations were negligible. The FWHM and the power index (2.8, see Sect. 2) of the Moffat function were fixed in the fitting process, while the PSF (NSC) central intensity was kept free and fitted.
@@
For the bulge and the disk we used respectively a Sérsic and a broken-exponential two-dimensional profiles (see Méndez-Abreu et al. 2017 for the specific implementation). We extensively tested different combinations of free and fixed parameters of these profiles, resulting in a better decomposition with a certain number of them fixed, as described below. Furthermore, the impact of small errors in the bulge and disk parameters on the NSC spectra, which is the goal of the decomposition, is negligible. The effective intensity, the effective radius and the Sérsic parameter of the bulge, as well as the central intensity of the disk, were kept free and fitted by the code, while the ellipticity and the position angle of both the bulge and the disk were fixed to the values from the S4G decomposition by Salo et al. (2015), who fitted the disk with a single exponential. We checked that the effective radius and the Sérsic parameter of the bulge resulting from the fit were very similar to the values from Salo et al. (2015). We did not expect to find the same values since they used a different wavelength range compared to ours.
@@
Before choosing a broken exponential for the disk profile, we performed a preliminary morphological analysis, fitting ellipses to the isophotes, in an image obtained from the entire PHANGS-MUSE mosaic, integrating all wavelengths in the analyzed spectral range. A one-dimensional radial profile, obtained by averaging azimuthally, follows a broken exponential, with the outer-profile scale length matching the disk scale length from S4G (Salo et al. 2015, using a single exponential). This is in agreement with the fact that the S4G team performed the morphological fit on a much larger spatial scale. We measured that the inner exponential has a scale length of 17.4 arcsec (${\sim}827$ pc) while the break radius is located at 30 arcsec (${\sim}1.4$ kpc) from the center. While our cropped data cube is dominated by the inner exponential, we fixed both these parameters in the C2D decomposition. With our new setup, GASP2D fitted each quasi-monochromatic image with a combination of a bulge, a disk and a NSC. This resulted in three data cubes, one for each component. Images resulting from integrating all wavelengths for each data cube are shown in Fig. 2. The light fractions of each component over the total are shown in two dimensions in Fig 3. This figure shows that the NSC dominates the light (with a light fraction above 60%) in the central spaxels within one PSF FWHM, while 100% of the light was assigned to the disk from a certain radius (${\sim}7$ arcsec) outwards.
Since we are interested in the stellar-population properties of the NSC and its comparison with the surrounding host galaxy, we will use for the rest of the analysis a data cube for the NSC and another for the rest of the galaxy. The latter was obtained by combining the bulge and the disk data cubes, and corresponds to the host galaxy, i.e. the galaxy after subtracting the NSC. We preferred this approach instead of considering separately the bulge and the disk, to minimize the impact of a potential suboptimal decomposition of these two morphological structures. This also follows previous studies of this kind, comparing the NSC to the host galaxy as a whole (Johnston et al. 2020; Fahrion et al. 2021). However, the C2D approach, providing spatially resolved spectral information, additionally allows us to fairly compare NSC properties with the host galaxy in the same aperture. This minimizes biases derived from quantitative comparisons of different spatial regions, due to spatial gradients in the host galaxy.
@@
We extracted integrated spectra in an aperture of radius $1\sigma$ of the PSF FWHM (${\sim}0.26$ arcsec $\approx 12$ pc) from these two data cubes and, for comparison, from the ”original” kinematic-corrected data cube before being processed by C2D (Sect. 4.1.1). Taking into account the central region of the PSF ensures that small errors in its estimate and variations of its size with wavelength have no impact on the extracted properties of the NSC. Hereafter, we will refer to: (1) the NSC spectrum and data cube as the one corresponding to the NSC component as extracted by C2D, (2) the ”host-galaxy” spectra and data cube as the one obtained combining the C2D disk and the bulge (and corresponding to the galaxy after subtracting the NSC), and (3) the ”original” spectrum and data cube as the original one before the C2D decomposition and containing the three morphological components (i.e., the full galaxy). These spectra are shown and compared to each other, for $1\sigma$ PSF, in Fig. 4.
[…]
@@ 4.2 Spectral fitting
We used a full-spectrum fitting technique both for the stellar population analysis (of all previously mentioned integrated spectra and data cubes), and for the calculation of the kinematic parameters necessary for their correction, as explained in Sect. 4.1.1 (step 3). We used the latest Python implementation of the penalized pixel-fitting (pPXF) method (Cappellari and Emsellem 2004; Cappellari 2017, 2023). pPXF models an observed spectrum as a linear combination of simple stellar population (SSP) templates, convolved with a line-of-sight velocity distribution (LOSVD) described by a Gauss–Hermite series. The resulting (light- or mass-fraction) weights, assigned to the SSP models to obtain the best-fit spectrum, can be used for the stellar-population analysis, while the LOSVD can be used to extract the kinematic parameters. In this work, different pPXF setups were used to extract stellar populations and kinematics. Stellar velocities and velocity dispersions were derived by fitting the central data cube resulting from step 2 in Sect. 4.1.1, with pPXF, using an additive polynomial of degree 8, and no multiplicative polynomial. Additive polynomials are typically used to minimize the effect of template mismatch in the kinematic fitting (Cappellari and Emsellem 2004; Cappellari 2017).
@@
We used in this work the semi-empirical sMILES SSP models (Knowles et al. 2023), the most up-to-date version of the MILES SSP models (Vazdekis et al. 2010, 2015). sMILES models are based on BaSTI isochrones (Pietrinferni et al. 2004, 2006) and offer five different values of the [Mg/Fe] abundances, covering the range between -0.2 and 0.6 dex (in steps of 0.2 dex). Throughout this paper, we will use [Mg/Fe] as a proxy for the enhancement in $\alpha$ elements. sMILES models cover ages between 0.03 and 14 Gyr, with 53 different values, and ten values of total metallicities [M/H] between -1.79 and 0.26 dex. This gives a total of 2650 SSP models. Since each of them corresponds to one combination of age, [M/H] and [Mg/Fe], coupling these models with pPXF allows us to extract the distribution of these three stellar-population parameters from each spectrum (given by the distribution of weights given to the different SSPs). These models cover a wavelength range between 3540 and 7409.6 $\mathrm{\SIUnitSymbolAngstrom}$, at a spectral resolution of 2.51 $\mathrm{\SIUnitSymbolAngstrom}$ at FWHM. Among sMILES sets of models, we chose the set of SSPs that adopted a bimodal initial mass function with a logarithmic slope of 1.3 for the segment of massive stars (Vazdekis et al. 2003). These models were normalized to correspond to one solar mass each, so that mass-weighted results are obtained from pPXF if SSPs are not renormalized. They need to be renormalized to their own median to obtain light-weighted results. We used in this work both normalizations, to obtain light-weighted stellar population maps and average results, and mass-weighted results to allow us to reconstruct SFHs in terms of mass fractions corresponding to different stellar populations.
@@
Our general approach for the stellar-population analysis was to average results from Monte Carlo simulations of unregularized pPXF fits, following Cappellari (2023). We used this approach to fit the three spectra extracted from the PSF aperture (Sect. 4.1.2), i.e. the NSC, the host-galaxy, and for comparison the spectrum from the original integrated data cube. We performed 100 Monte Carlo realizations for each spectrum. For each realization, we perturbed the spectrum with random, wavelength-dependent noise drawn from a Gaussian distribution. The noise level was taken from the residuals at the specific wavelength, from an initial (unregularized) pPXF fit. From the distribution of SSP weights for the 100 realizations, we estimated the mean ages, [M/H] and [Mg/Fe], and their statistical uncertainties (as standard deviations), for each of the three original spectra. This allowed us not to use any regularization on the SSP weights during the pPXF fits.
@@
The resulting host-galaxy data cube from the decomposition was rebinned to a higher S/N of 100 before performing the spectral fitting for the spatially resolved stellar population analysis, ensuring robust results and lower uncertainties in the outer fainter region. This still left us with 6766 Voronoi bins, which made the Monte Carlo approach time-consuming and computationally expensive for the entire data cube. Thus, we decided to map the stellar-population parameters of the host galaxy using pPXF regularized fits with a unit regularization parameter. While one single unregularized fit gives discrete and noisy solutions, some level of regularization ensures smoother and more realistic stellar-population solutions (e.g., Cappellari 2017). However, regularization above a certain level can lead to biases in the results and dilute features in the SFHs. Thus, we used the approach presented in Pinna et al. (2019b) to choose the regularization parameter (described as follows, see also Pinna et al. 2019a; Martig et al. 2021; Sattler et al. 2023, 2025). After rescaling the pPXF noise parameter as the one necessary to obtain a reduced $\chi^{2}=1$ with no regularization, we choose a regularization well below a certain maximum calibration value. We defined the calibration regularization as the minimum value increasing $\chi^{2}$, from the case with no regularization, by $\Delta\chi^{2}\approx\sqrt{2N_{\text{goodpix}}}$, where $N_{\text{goodpix}}$ is the number of fitted (spectral) pixels (McDermid et al. 2015; Cappellari 2017; Boecker et al. 2020). For comparison, we also mapped the stellar populations from the original data cube before the C2D decomposition, using this same approach (S/N=100=100, multiplicative polynomial of degree 8 and unit regularization parameter).